i cant find the surface area of the triangular prism The base of the prism is an isosceles triangle.
the numbers are 24 as a base, 35 as a height, 46 as a length then it has 37 on the side

Answer by rothauserc(4718)

surface area of triangular prism(SA) = (2 * base triangle's area) + (triangle's perimeter * prism's height)
:
Note that base triangle's area = 2 * ((1/2 * base * height) = base * height
:
base triangle's area = 24 * 35 = 840
:
base triangle's perimeter = 24 + (2 * 37) = 98
:
***************************
SA = 840 + (98 * 46) = 5348
***************************
: